OpenStudy (anonymous):

Simplify 18 times y to the third power all over 36 times x times y squared. Answer y squared over 2 times x 1 over 2 times x times y 1 over 2 times x y over 2 times x

OpenStudy (anonymous):

18y^3/36xy^2 = 18y/36x = y/2x

OpenStudy (anonymous):

y over 2 times x 18 times y to the third power all over 36 times x times y squared. 18y³/ 36x y² = y/ 2x